OpenStudy (anonymous):

How do i solve this s = π * r * √ (r^2 + h^2)

OpenStudy (anonymous):

for what variable? r or h?

OpenStudy (anonymous):

r

OpenStudy (anonymous):

\[r = {-{\sqrt{-{\sqrt{\pi^2h^4+4s^2} \over \pi} -h^2} \over \sqrt 2}}\]